use factoring to solve the polynomial equation
8x^4 -72x^2 = 0

Answer by LinnW(1048)

8x^4 -72x^2 = 0
8x^2(x^2 - 9) = 0
8x^2(x + 1)(x - 1) = 0
So x = 0 or -1 or 1